Use a double integral in polar coordinates to find the area of the region common to the interior of the cardioids

r = 1 + cos θ and r = 1 − cos θ.


1
Expert's answer
2021-11-22T17:57:36-0500


In order to calculate the area between the curves, we determine the points of intersection:


"1+\cos \theta=1-\cos \theta"

"\cos \theta=0"

"\theta=\pi/2"

"\theta=-\pi/2"

The area is calculated using the following equation:


"A=\displaystyle\int_{-\pi/2}^{\pi/2}\displaystyle\int_{1-\cos \theta}^{1+\cos \theta}rdrd\theta"

Using symmetry:


"A=\displaystyle\int_{-\pi/2}^{\pi/2}\displaystyle\int_{1-\cos \theta}^{1+\cos \theta}rdrd\theta=2\displaystyle\int_{0}^{\pi/2}\displaystyle\int_{1-\cos \theta}^{1+\cos \theta}rdrd\theta"

"=2\displaystyle\int_{0}^{\pi/2}\bigg[\dfrac{r^2}{2}\bigg]\begin{matrix}
 1+\cos \theta \\
 1-\cos\theta
\end{matrix}d\theta"

"=2\displaystyle\int_{0}^{\pi/2}(4\cos\theta)d\theta=8[\sin \theta]\begin{matrix}
 \pi/2 \\
 0
\end{matrix}=8({units}^2)"